ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 104, 1998

Bahturin, Y., A. Giambruno and D. M. Riley Group-graded algebras with polynomial identity

145

Ballico, E. and C. Keem Weierstrass multiple points on algebraic curves and ramified coverings

335

Barnea, Y., A. Shalev and E. I. Zelmanov Graded subalgebras of affine Kac–Moody algebras

321

Donin, J. and S. Shnider Deformation of certain quadratic algebras and the corresponding quantum semigroups

285

Giambruno, A. See Y. Bahturin Ginzburg, D. and Z. Rudnick Stable multiplicities in the length

spectrum of Riemann surfaces

129

Hájek, P. Smooth functions on c_{0}  17

Hernández, F. L. and B. Rodriguez-Salinas Lattice-embedding scales of L^{p} spaces into Orlicz spaces

191

del Junco, A. A simple map with no prime factors  301

Keem, C. See E. Ballico Neumann, K. Every finitely generated regular field extension has a

stable transcendence base

221

Pucci, P. and J. Serrin Local asymptotic stability for dissipative wave systems

29

Riley, D. M. See Y. Bahturin Rodriguez-Salinas, B. See F. L. Hernández Rudnick, Z. See D. Ginzburg

de la Rue, T. Rang des systèmes dynamiques gaussiens  261

Serrin, J. See P. Pucci Shalev, A. See Y. Barnea Shnider, S. See J. Donin

Sparks, N. A distortion problem for the space ℓ_{∞}^{c}(ω_{1})  125

Steel, J. R. and P. D. Welch Σ_{3}^{1} absoluteness and the second uniform indiscernible

157

Terenzi, P. A positive answer to the basis problem  51
Valtr, P. Guarding galleries where no point sees a small area  1

Welch, P. D. See J. R. Steel Zelmanov, E. I. See Y. Barnea